# Low-Rank Continual Personalization of Diffusion Models

### v1

@@ Abstract
Recent personalization methods for diffusion models, such as Dreambooth, allow fine-tuning pre-trained models to generate new concepts. However, applying these techniques across multiple tasks in order to include, e.g., several new objects or styles, leads to mutual interference between their adapters. While recent studies attempt to mitigate this issue by combining trained adapters across tasks after fine-tuning, we adopt a more rigorous regime and investigate the personalization of large diffusion models under a continual learning scenario, where such interference leads to catastrophic forgetting of previous knowledge. To that end, we evaluate the naïve continual fine-tuning of customized models and compare this approach with three methods for consecutive adapters’ training: sequentially merging new adapters, merging orthogonally initialized adapters, and updating only relevant parameters according to the task. In our experiments, we show that the proposed approaches mitigate forgetting when compared to the naïve approach.
11footnotetext: Equal contribution.
@@ 1 Introduction
Deep Diffusion Probabilistic Models (DDPMs) [28] have revolutionized image generation, with state-of-the-art performance in synthesizing highly detailed objects and blending them with complex styles. Moreover, personalization techniques such as Dreambooth [24], combined with the Low-Rank Adaptation (LoRA) [12], introduce the possibility to easily customize large-scale pre-trained models. In those techniques, a selected subset of model parameters is updated with just a few training examples to allow for the generation of objects not present in the original training set.
Unfortunately, continuous customization of a generative model over several tasks leads to the catastrophic forgetting [6] of previously encoded knowledge due to mutual interference between tasks’ adapters. Recent studies try to mitigate this issue through additional regularization [27] or by careful merging of the trained adapters [26, 7]. For such merging to succeed, all adapters have to be available at the same time to properly scale [26] or select [18] the resulting models’ weights. In practice, with a bigger number of tasks, this assumption is often unfeasible, as storing numerous copies of model weights, even in the form of a sparse low-rank adaptation matrix, is memory-consuming (e.g. approximately 8 LoRA weight matrices with rank $64$ are equal to the size of all adapted parameters in Stable Diffusion XL’s [21] model). Therefore, in this work, we propose to study the effectiveness of merging techniques under the strict continual learning regime where only the model or model with a single adapter are passed between tasks.
In a series of experiments, we evaluate (1) Naïve continual fine-tuning of the low-rank adapter with three merging methods mitigating forgetting of concepts and styles: (2) consecutive merging of task-specific adapters, (3) merging LoRAs initialized with orthogonal weights, and (4) merging through a selection of weights with the highest magnitude for the task [18]. Our experiments indicate that adding multiple adapters in a Naïve way can lead to a situation where a model converges towards its base form, while all the evaluated techniques mitigate this issue.
@@ 2 Related work
The topic of Diffusion Models (DMs) personalization has been broadly studied in recent years. Several works, including Textual Inversion [8] learn new word embeddings to represent custom and unique concepts. This idea is further extended by several approaches [17, 16], with additional regularization [33], or features mapping [32]. On the other hand, DreamBooth [24] fine-tunes the entire model to incorporate specific subjects into generated images. This idea is further extended in HyperDreamBooth [25] where a hypernetwork [10] is used to generate personalized weights. Recent personalization techniques employ Low-Rank Adaptation (LoRA) [12] to reduce the computational requirements by fine-tuning small, low-rank matrices and keeping the base model frozen. Nevertheless, simply using LoRA does not prevent interference. Therefore, ZipLoRA [26] optimizes separately trained adapters for style and object to simplify their merging, and B-LoRA [7] targets different attention layers of diffusion models for both style and object adapters. On top of those, weights2weights [5] introduces a personalized adapter space, enabling the creation of entirely new personalized models. OFT [22] and Orthogonal Adaptation [20] leverage orthogonalization to merge models seamlessly without compromising fidelity.
@@
Apart from introducing new objects to the model, there is a line of research aiming for style generation and editing in DDPMs. StyleDrop [29] leverages the Muse transformer to fine-tune text-to-image models for generating style-specific images. StyleAligned [11] employs shared attention to maintain a consistent style across generated images, while Style Similarity [30] provides a framework for style retrieval from diffusion models’ training data and introduces CSD, an image encoder that accurately captures the style of images. UnlearnCanvas [36] offers a comprehensive dataset for evaluating machine unlearning in stylized image generation. For knowledge localization, Diff-QuickFix [2] and LOCOEDIT [1] enable precise editing of diffusion models’ knowledge of specific styles by identifying style-relevant layers.
@@
There are works using DMs in CL scenarios. In [35], the authors explore if contemporary CL methods prevent forgetting in diffusion models. In [9], DDPM is used as a source of rehearsal examples, while [4] extends this idea with classifier guidance towards easy-to-forget examples. Similarly, in [15], authors use an externally trained text-to-image diffusion model for the same purpose. To mitigate artifacts introduced with generative replay, especially with low-rank adaptation, in [27], the authors introduce C-LoRA, where consecutive tasks are learned with a self-regularization technique that penalizes updates corresponding to already altered weights.
@@
The concept of editing models through arithmetic operations applied on models’ weights is explored in [13], where authors describe mathematical operations between task vectors, defined as the difference between the weights of the fine-tuned and the base model. To limit the interference between tasks, recent methods replace straightforward averaging with summation of trimmed parameters [34], Fisher-information-based merging [19] or with guidance from discriminative model [14]. In this work, we adapt the MagMax method [18] that is used to continually train the CLIP model by selecting, from several task vectors, weights with the highest magnitude.
@@ 3 Methods
We evaluate four different approaches to continual model personalization. Notation-wise, we denote $W_{0}$ as the base model weights, while $(A_{t},B_{t})$ represent the fine-tuned low-rank adapter resulting from personalizing DM on the $t$-th task. New adapter weights $(A_{t},B_{t})$ are calculated through fine-tuning on data from task $t$, given the frozen base model weights $W_{0}$, and an adapter fine-tuned over $t-1$ tasks denoted as $(A_{1..t-1},B_{1..t-1})$. The final model can be calculated as $W_{1...T}=W_{0}+B_{1...T}A_{1...T}$. All the evaluated strategies differ in initialization or merging strategy, which we overview in Fig. 1.
Figure 1: Comparison of initialization and merging in four evaluated methods. The expressions in the adapters refer to how they are initialized with a new task.
@@
In this naïve approach, we sequentially fine-tune the adapter weights from the previous task: $A_{t-1}$, $B_{t-1}$ without reinitializing them. The intermediate adapters are not merged into the base model.
@@
In the second approach, we create a new LoRA adapter for each task with a standard initialization technique ($A\sim N(0,I),B=0$). Then, at the end of each task, we merge the adapter with the base model and pass it to the next task.
@@
We further extend the merging and reinitialization approach and limit the interference between tasks by pre-directing the fine-tuning process toward non-conflicting adapters by initializing $B_{t}$ weights with zeros and setting $A_{t}$ weights as orthogonal to $A_{1..t-1}$ using Singular Value Decomposition (SVD). We decompose the $i$-th column of $A_{1..t-1}$ as:
${A^{(i)}_{1..t-1}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$ (1)
and select the last row of the conjugate transpose of the right singular vectors $\mathbf{V}$, corresponding to the smallest singular value.
(4) Magnitude-based selection of LoRA weights
In the last approach, we adapt the magnitude-based parameter selection introduced in [18]. To that end, after each task, we select parameters with the highest magnitude when comparing already merged adapters $(A^{\prime}_{1..t-1},B^{\prime}_{1..t-1})$ with the current one $(A_{t},B_{t})$. The resulting merged adapter is passed together with a base model to the new task. Task-related adapters are always initialized in the standard way, while for the frozen Base model, we use a version with parameters updated by the merged adapter.
@@ 4 Experiments
Experimental setup
In our experiments, we use the DreamBooth [24] dataset for fine-tuning on objects, and UnlearnCanvas [36] for styles. We perform fine-tuning on the $10$ consecutive tasks and average results over 4 random tasks ordering, with 2 random seeds each. To measure the performance of the model, we use cosine similarity to the reference image with DINO [3] and CLIP [23] feature extractors for object alignment measure and DINO and CSD [31] for style similarity. We report the performance of the final model fine-tuned over all 10 tasks with the Average Score and Average Forgetting metrics. More details about the experimental setup are provided in the Appendix A.
Continual Personalization Methods Comparison
In Tab.1 and Tab.2, we present the results for the continual object and style personalization, respectively. For both tasks, we observe that the merge & initialization method, either with or without orthogonalization, outperforms naïve continual fine-tuning. On the other hand, magnitude-based merging has the lowest average forgetting at the expense of lower general performance as a result of the lower adaptation capabilities of the method. While magnitude-based weights selection yields state-of-the-art performance in discriminative tasks [18], selecting a limited number of weights hinders the precise merging of new generative traits. Detailed analysis in Appendix B on performance after each task shows that the merge & orthogonalization approach exhibits the best plasticity at the expense of the highest forgetting.
@@
Merge & Initialization $\boldsymbol{0.675}_{\pm.005}$ $\boldsymbol{0.457}_{\pm.011}$ $\underline{0.026}_{\pm.003}$ $\underline{0.056}_{\pm.009}$
@@
Magnitude-based Merging $0.643_{\pm.002}$ $\underline{0.408}_{\pm.006}$ $\boldsymbol{0.018}_{\pm.002}$ $\boldsymbol{0.036}_{\pm.006}$
Table 1: Average Score and Average Forgetting in CLIP-I and DINO object alignment metrics for our continual personalization methods. The best results are in bold, while the second best are underlined.
@@
Merge & Initialization $\boldsymbol{0.385}_{\pm.019}$ $\boldsymbol{0.285}_{\pm.019}$ $\underline{0.131}_{\pm.019}$ $\underline{0.085}_{\pm.021}$
@@
Magnitude-based Merging $0.289_{\pm.015}$ $0.240_{\pm.009}$ $\boldsymbol{0.093}_{\pm.013}$ $\boldsymbol{0.052}_{\pm.015}$
Table 2: Average Score and Average Forgetting in CSD and DINO style alignment metrics for our continual personalization methods. The best results are in bold, while the second best are underlined.
To further analyze the results, in Fig 2 (left), we present a comparison of different methods’ performance on the first task after continual adaptation on the given number of tasks. We can observe that the naïve approach decays to the performance of the baseline model, while other techniques can mitigate forgetting better. In Fig 2 (right), we present the possible reason for adapters’ degradation.
We measure the extent of the mutual interference between tasks during training. To that end, we create a vector build of directions that oppose the adapters of a task after training on subsequent tasks and plot the ratio of this vector’s norm to the original task vector norm. An explanation of calculating the degradation of adapters, along with the detailed results of our evaluation for the style adaptation task, are presented in Appendix B.
Refer to caption Refer to caption
Figure 2: DINO score on the first task over continual fine-tuning on next tasks for the object customization (left) and the ratio of opposed parameters vector norm to the task adapter norm (right).

### v3

@@ Abstract
Recent personalization methods for diffusion models, such as Dreambooth and LoRA, allow fine-tuning pre-trained models to generate new concepts. However, applying these techniques across consecutive tasks in order to include, e.g., new objects or styles, leads to a forgetting of previous knowledge due to mutual interference between their adapters. In this work, we tackle the problem of continual customization under a rigorous regime with no access to past tasks’ adapters. In such a scenario, we investigate how different adapters’ initialization and merging methods can improve the quality of the final model. To that end, we evaluate the naïve continual fine-tuning of customized models and compare this approach with three methods for consecutive adapters’ training: sequentially merging new adapters, merging orthogonally initialized adapters, and updating only relevant task-specific weights. In our experiments, we show that the proposed techniques mitigate forgetting when compared to the naïve approach. In our studies, we show different traits of selected techniques and their effect on the plasticity and stability of the continually adapted model. Repository with the code is available at https://github.com/luk-st/continual-lora.
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015) (DMs) have revolutionized image generation with state-of-the-art performance in synthesizing detailed objects and blending them with complex styles. Personalization techniques such as Dreambooth (Ruiz et al., 2023), combined with the Low-Rank Adaptation (LoRA, Hu et al. (2021)), introduce the possibility to easily customize large-scale pre-trained models. In those techniques, a selected subset of model parameters is updated with just a few training examples to enable the generation of concepts not present in the original training set.
Unfortunately, customization of a generative model over several tasks leads to the catastrophic forgetting (French, 1999) of previous knowledge due to mutual interference between tasks’ adapters. Recent studies (Shah et al., 2023; Frenkel et al., 2024) try to tackle this issue in terms of object and style customization tasks’ adapters by carefully merging the trained adapters, either with optimization (Shah et al., 2023) or by targeting different attention layers (Frenkel et al., 2024). However, for the merging to succeed, the adapters have to be available at the same time to properly scale (Shah et al., 2023) or select (Marczak et al., 2024) the resulting models’ weights. Even though LoRA reduces the number of adapted parameters, in practice, with many objects or styles, it is unfeasible to store copies of model weights for each of them. Even in the form of sparse matrices, keeping all the adapters is memory-consuming (e.g., approximately 8 LoRA weight matrices with rank $64$ are equal to the size of all adapted parameters in Stable Diffusion XL’s (Podell et al., 2024) model).
Therefore, in this work, we propose to study the effectiveness of LoRA weights initialization and merging under the strict continual learning regime where only the model or model with a single adapter is passed between tasks. In a series of experiments, we compare (1) Naïve continual fine-tuning of the low-rank adapter, and three simple merging baselines mitigating forgetting of concepts and styles: (2) consecutive merging of task-specific adapters, (3) merging LoRAs initialized with orthogonal weights, and (4) merging through a selection of weights with the highest magnitude for the task. Our experiments indicate that adding multiple adapters in a Naïve way can lead to a situation where a model converges, in its performance, towards its base form, while all the evaluated techniques mitigate this issue.
We show the significant distinctions between different initialization methods. For example, orthogonalized initialization results in high plasticity but low stability, whereas selection-based merging exhibits the opposite trend. Our findings indicate that optimal performance is achieved by reinitializing the LoRA weights and merging them using a standard approach. We hypothesize that this effect arises from the specific initialization of LoRA weights, which facilitates sparse adaptation of the base model weights.
@@ 2 Related work
The topic of Diffusion Models personalization has been broadly studied in recent years. Several works, including Textual Inversion (Gal et al., 2022) learn new word embeddings to represent custom and unique concepts. This idea is further extended by Li et al. (2023); Kumari et al. (2023), with additional regularization (Wu et al., 2024), or features mapping (Wei et al., 2023). On the other hand, DreamBooth (Ruiz et al., 2023) fine-tunes the entire model to incorporate specific subjects into generated images. This idea is further extended in HyperDreamBooth (Ruiz et al., 2024) where a hypernetwork (Ha et al., 2016) is used to generate personalized weights. Recent personalization techniques employ Low-Rank Adaptation to reduce the computational requirements by fine-tuning small, low-rank matrices and keeping the base model frozen. Nevertheless, the simple merging of weights fine-tuned with LoRA does not prevent task interference. Therefore, in ZipLoRA Shah et al. (2023) optimize separately trained adapters for style and object to simplify their merging, and Frenkel et al. (2024) (B-LoRA) target different attention layers of diffusion models for both style and object adapters. OFT (Qiu et al., 2023) and Orthogonal Adaptation (Po et al., 2024) leverage orthogonalization to merge models seamlessly without compromising fidelity. Finally, Dravid et al. (2024) introduce weights2weights – a personalized adapter space, enabling the creation of entirely new personalized models.
@@
Apart from introducing new objects to the model, there is a line of research aiming for style generation and editing in Diffusion Models. StyleDrop (Sohn et al., 2023) leverages the Muse transformer to fine-tune text-to-image models for generating style-specific images. StyleAligned (Hertz et al., 2023) employs shared attention to maintain a consistent style across generated images, while Style Similarity (Somepalli et al., 2024b) provides a framework for style retrieval from diffusion models’ training data and introduces CSD, an image encoder that accurately captures the style of images. UnlearnCanvas (Zhang et al., 2024) offers a comprehensive dataset for evaluating machine unlearning in stylized image generation. Diff-QuickFix (Basu et al., 2024b) and LOCOEDIT (Basu et al., 2024a) enable precise editing of diffusion models’ knowledge of specific styles by identifying style-relevant layers.
@@
There is a growing line of research employing DMs in continual learning (CL) mainly as a source of rehearsal examples from past tasks (Gao & Liu, 2023). This idea is extended by Cywiński et al. (2024) through classifier guidance towards easy-to-forget examples. Similarly, Jodelet et al. (2023) uses an externally trained text-to-image diffusion model for the same purpose. On the other hand, Zając et al. (2023) explore if contemporary CL methods prevent forgetting in diffusion models.
Smith et al. (2023) introduce the task of continual customization, with the goal of fine-tuning the diffusion model on consecutive personalization tasks. They introduce C-LoRA, where consecutive tasks are learned with a self-regularization technique that penalizes updates corresponding to the already altered weights. STAMINA (Smith et al., 2024) introduces learnable adapter’s attention masks parameterized with low-rank MLPs for better scaling to longer task sequences. CIDM (Dong et al., 2024) develop a concept-consolidation loss that explores both the distinctive characteristics of concepts within each task (Task-Specific Knowledge) and shared information across different tasks with similar concepts (Task-Shared Knowledge). Finally, Jha et al. (2024) propose to leverage classifier scores into both parameter-space (to improve the task-specific Fisher information estimates of EWC) and function-space regularizations (creating a double-distillation framework), achieving high performance in continual personalization. While prior methods improve knowledge consolidation using additional regularizations, in this work, we study the fundamental properties of the most basic components for continual personalization which are merging and initialization techniques.
@@
The concept of editing models through arithmetic operations applied on models’ weights is explored by Ilharco et al. (2023), where authors describe mathematical operations between task vectors, defined as the difference between the weights of the fine-tuned and the base model. To limit the interference between tasks, recent methods replace straightforward averaging with summation of trimmed parameters (Yadav et al., 2024), Fisher-information-based merging (Matena & Raffel, 2022) or with guidance from discriminative model (Jin et al., 2022). In this work, we adapt MagMax (Marczak et al., 2024) to our experimental setting, the method achieving state-of-the-art performance in both class-incremental and domain-incremental learning in discriminative modeling. MagMax sequentially fine-tunes the model and consolidates the knowledge thanks to maximum magnitude weight selection from several low-rank task adapters.
@@ 3 Methods
We evaluate four different approaches to continual diffusion model personalization. Notation-wise, we denote $W_{0}$ as the base model weights, while $\{A_{t},B_{t}\}$ represent the fine-tuned low-rank adapter resulting from personalizing DM on the $t$-th task. New adapter weights $\{A_{t},B_{t}\}$ are calculated through fine-tuning on data from task $t$, given the frozen base model weights $W_{0}$, and an adapter fine-tuned over $t-1$ tasks denoted as $\{A_{1\dots t-1},B_{1\dots t-1}\}$. The final model can be calculated as $W_{1\dots T}=W_{0}+B_{1\dots T}A_{1\dots T}$. All the evaluated strategies differ in either initialization or merging strategy, which we overview in Fig. 1, and describe further in this section.
Figure 1: Comparison of initialization and merging in four evaluated methods. The expressions in the adapters refer to how they are initialized with a new task. Numbers next to each schema refer to textual descriptions in Section 3.
@@
In the naïve approach, we sequentially fine-tune the adapter weights from the previous task $\{A_{t-1},B_{t-1}\}$ with the current task data without any reinitialization. Thus, the intermediate adapters, which are continually changed, can be merged into the base model after completing all the $T$ tasks.
@@
We create a new LoRA adapter for each task with a standard initialization technique ($A\sim\mathcal{N}(0,\mathcal{I}),B=0$). Then, at the end of each task, we merge the adapter with the base model and pass it to the next task. This approach is similar to naïve finetuning. However, thanks to the reinitialization of the LoRA parameters with random values for the $A$ matrix, each task is optimized in its own independent low-dimensional space. Therefore, compared to a naïve approach, the Merge & Initialization technique naturally reduces task interference, as each adapter independently learns only a minimal set of task-specific parameters. Marczak et al. (2025) show that such situation has positive impact on the final model perfromance in discriminative modeling. Intuitively, this can be related to the two-stage training (Kamra et al., 2017) introduced for various CL methods.
@@
We further extend the merging and reinitialization approach in order to limit the interference between tasks by initializing, for the given task $t$, $B_{t}$ weights with zeros and setting $A_{t}$ weights’ columns as orthogonal to columns of $A_{1\dots t-1}$ matrix using Singular Value Decomposition (SVD). We decompose the $i$-th column of $A_{1\dots t-1}=\Sigma_{s\in 1\dots t-1}A_{s}$ as
${A^{(i)}_{1\dots t-1}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}\vspace{-0.3em}$ (1)
and select the last row of the conjugate transpose of the right singular vectors $\mathbf{V}$, corresponding to the smallest singular value. The underlying assumption of this approach is to pre-direct the fine-tuning process toward the non-conflicting adapters, thus limiting the mutual drift of weights that are introduced by consecutive tasks.
(4) Magnitude-based selection of LoRA weights.
Finally, we adapt and modify the magnitude-based parameter selection method introduced in Marczak et al. (2024). As the original MagMax approach involves merging adapters into the model after all the $T$ tasks, continuing the method for $(T+1)$-th task is impossible without storing base model weights $\theta_{0}$, weights obtained after $T$ tasks $\theta_{T}$ (before selection operation), and the running max-magnitude statistics $\tau_{\textsc{M{ag}M{ax}}}$. To that end, after each task $t$, we run the MagMax selection of parameters with the highest magnitude by comparing the sum of already merged adapters $\{A_{1\dots t-1},B_{1\dots t-1}\}$ with the current one $\{A_{t},B_{t}\}$:
$\{A_{1\dots t},B_{1\dots t}\}=\{\textsc{M{ag}M{ax}}(A_{1\dots t-1};A_{t}),\textsc{M{ag}M{ax}}(B_{1\dots t-1};B_{t})\}.\vspace{-0.1em}$ (2)
The resulting adapter is passed together with a base model into the next task. At the beginning of a new task, the so far selected adapters are temporarily merged into the frozen Base model, while the new adapters’ parameters are initialized in the standard way, optimized, and passed to the selection procedure (2).
@@ 4 Experiments
Experimental setup.
In our experiments, we use the DreamBooth (Ruiz et al., 2023) dataset for fine-tuning on objects, and UnlearnCanvas (Zhang et al., 2024) for styles. We perform fine-tuning on the $10$ consecutive tasks and average results over 4 random tasks ordering, with 2 random seeds each. To measure the performance of the model, we calculate cosine similarity between model outputs and reference images with DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) feature extractors for object alignment measure and DINO and CSD (Somepalli et al., 2024a) for style similarity. We report the performance of the final model fine-tuned over all 10 tasks with the Average Score and Average Forgetting metrics. More details about the experimental setup are provided in the Appendix A.
Results.
In Table 1 and Table 2, we present the results for the continual object and style personalization, respectively. For both tasks, we observe that the merge & initialization method, either with or without orthogonalization, outperforms naïve continual fine-tuning. On the other hand, magnitude-based merging has the lowest average forgetting at the expense of lower general performance as a result of the lower adaptation capabilities of the method. While magnitude-based weights selection yields state-of-the-art performance in discriminative tasks (Marczak et al., 2024), selecting a limited number of weights hinders the precise merging of new generative traits. Detailed analysis in Appendix B on performance after each task shows that the merge & orthogonalization approach exhibits the best plasticity. However, a simple summation of orthogonal task vectors makes Merge & Orthogonalization fail in retaining original performance on past tasks, leading to the highest forgetting. The Merge & Initialization method balances well a relatively high plasticity with stability, as observed by well-preserved past knowledge.
@@ Results.
Merge & Initialization $\bm{0.675}_{\pm.005}$ $\bm{0.457}_{\pm.011}$ $\underline{0.026}_{\pm.003}$ $\underline{0.056}_{\pm.009}$
@@
Magnitude-based Merging $0.643_{\pm.002}$ $\underline{0.408}_{\pm.006}$ $\bm{0.018}_{\pm.002}$ $\bm{0.036}_{\pm.006}$
Table 1: Average Score and Average Forgetting in CLIP-I and DINO alignment metrics for continual object personalization. The best results are in bold, while the second best are underlined.
@@
Merge & Initialization $\bm{0.385}_{\pm.019}$ $\bm{0.285}_{\pm.019}$ $\underline{0.131}_{\pm.019}$ $\underline{0.085}_{\pm.021}$
@@
Magnitude-based Merging $0.289_{\pm.015}$ $0.240_{\pm.009}$ $\bm{0.093}_{\pm.013}$ $\bm{0.052}_{\pm.015}$
Table 2: Average Score and Average Forgetting in CSD and DINO alignment metrics for continual style personalization. The best results are in bold, while the second best are underlined.
In Fig. LABEL:fig:task1_and_ratio_objects, we compare the performance of different methods on the first task after continual adaptation across multiple tasks in object (left) and style (right) personalization. The results show that the naïve approach degrades overall quality to the baseline model’s performance, while other techniques mitigate forgetting more effectively. Notably, knowledge retention varies significantly between object and style personalization. While the Merge & Initialization method reduces forgetting in both cases, resetting weights with orthogonal initialization performs significantly worse in style personalization. We hypothesize that this is due to task vectors being more aligned in style personalization. In such cases, more orthogonal vectors lead to higher forgetting.